The admission fee at the steamboat springs hot air balloon festival is 1.50 for children and 4.00 for adults. On a certain day 1130 people enter the festival and 2520 is collected. How many children and adults attended? cant figure out a systemically equation for this been stuck for an hour.
Answer by jorel555(1290)

Let m be the number of children at the festival, and n be adults. Then:
m+n=1130
and
1.5m+4n=2520
So:
4m+4n=4520
1.5m+4n=2520
2.5m=2000
m=800
n=330
There were 800 children and 330 adults at the festival. ☺☺☺☺